# Multi-Channel Hypergraph Contrastive Learning for Matrix Completion

### v1

@@ 3. Problem Definition
We set $\mathbf{U}=\{u_{1},u_{2},\ldots,u_{M}\}$ as the set of users and $\mathbf{V}=\{v_{1},v_{2},\ldots,v_{N}\}$ as the set of items, where $M$ and $N$ are the number of users and items, respectively. We use a binary matrix $\mathbf{R}$ to store historical ratings, where $r_{uv}$ indicates a rating from user $u$ to item $v$, and $r_{uv}=0$ means item $v$ is unexposed to user $u$. Interaction data is represented as a user-item bipartite graph $G$, and the task is to predict missing ratings in this graph.
Recent developments in graph neural networks have predominantly applied matrix completion techniques to bipartite user-item graphs, treating it as a link prediction problem. This approach utilizes GNNs to capture complex relationships and dependencies within user-item interactions, advancing recommender systems and collaborative filtering approaches.
Key notations are summarized in Table 5 in Supplement.
@@ 4. Methodology
In this section, we provide a comprehensive overview of the proposed MHCL framework, as illustrated in Figure 1. MHCL comprises four key components: (1) Subgraph Partition and Graph Convolution: MHCL begins by partitioning the bipartite user-item graph into subgraphs based on different ratings. Graph convolutions are applied to these subgraphs to capture collaborative signals and extract local topological features for users and items across various rating contexts. (2) Multi-Channel Cross-Rating Contrastive Learning: To enhance interaction between adjacent rating subgraphs, MHCL employs a multi-channel cross-rating contrastive learning module. This module promotes information exchange between rating channels, addressing data sparsity and mitigating the long-tail distribution problem. (3) Adaptive Hypergraph Structure Learning: Using local embeddings from the subgraphs, MHCL integrates an adaptive hypergraph structure learning module. This module constructs hypergraphs by identifying high-order dependencies among nodes of the same type (e.g., users with users, items with items), capturing global topological features unique to each rating. This process helps mitigate over-smoothing and enhances the model’s ability to learn higher-order correlations. (4) Attention-Based Cross-View Aggregation: MHCL aggregates local and global representations through an attention-based cross-view aggregation module. By maximizing the consistency between embeddings from diverse views, it leverages global-local contrastive learning techniques to obtain robust final representations.
Through these four components, MHCL delivers a comprehensive solution for matrix completion, leveraging collaborative signals, topological features, and higher-order dependencies to enhance the accuracy and efficiency of recommendations.
@@ 4.1. Subgraph Partition and Graph Convolution
(1) $x_{r}^{0}[i]=Concat(\mathbf{E}_{r}[\mathbf{id}[i]],\mathbf{E}_{r}[\mathbf{ra}[i]],\mathbf{E}_{r}[\mathbf{oh}[i]]),$
where $\mathbf{E}_{r}\in\mathbb{R}^{(M+N+3)\times d}$ denotes $r$-th trainable parameter matrix in $\mathbf{E}\in\mathbb{R}^{\mathcal{|R|}\times(M+N+3)\times d}$, and $x_{r}^{0}[i]$ denotes the initializing embedding of the $i$-th node in the $r$-th rating subgraph. The node embeddings are the input of the message-passing layer.
Then we systematically partition the graph into distinct rating subgraphs based on the different types of rating edges. For example, with five rating types, we create five rating subgraphs, each containing node embeddings specific to its rating category. MHCL then operates independently on these subgraphs, applying graph convolutions to derive node embeddings within each subgraph. Unlike conventional GCNs, MHCL forgoes the use of a learnable matrix for feature transformation, opting for an explicit propagation process, which can be expressed as follows:
@@ 4.2. Multi-Channel Cross-Rating Contrastive Learning
Unlike traditional heterogeneous graphs with multiple edge types (e.g., click, purchase), our scenario focuses on comparable ratings, where adjacent ratings indicate similar user preferences and node representations. However, the long-tail distribution in real-world datasets complicates MHCL’s partitioning into distinct subgraphs based on rating types, risking the loss of magnitude and order information during graph convolution.
To address this, we introduce a multi-channel cross-rating contrastive learning module. Each rating subgraph acts as a separate channel, capturing unique perspectives, while cross-rating contrastive learning identifies cooperative associations. By using the InfoNCE loss (Gutmann and Hyvärinen, 2010), our approach improves model performance by treating embeddings from adjacent rating subgraphs as positive pairs, preserving the rating order:
(3) ${\mathcal{L}_{S(k)}^{\mathbf{U}}=\sum_{u\in\mathbf{U}}-\log\frac{exp((x_{u}^{(k)}\cdot x_{u}^{(k+1)}/\tau))}{\sum_{u^{\prime}\in\mathbf{U}}exp((x_{u^{\prime}}^{(k)}\cdot x_{u^{\prime}}^{(k+1)}/\tau))},}$
where $x_{u}^{(k)}$ is the normalized output embedding of $k$-th rating subgraph and $\tau$ is the temperature hyper-parameter of softmax.
@@ 4.3. Adaptive Hypergraph Structure Learning
The embedding from Eq. (2) captures collaborative signals and topological information within rating subgraphs, reflecting implicit user preferences or item features. To address the limitations of bipartite graphs, where edges connect nodes of the same type, MHCL employs hypergraph structure learning for users and items. By utilizing hypergraph convolution, MHCL captures global information beyond node types. Inspired by previous work (Wei et al., 2022; Xia et al., 2022), we use an MLP to compute hyperedge assignments within each rating subgraph, resulting in hypergraphs $\mathbf{H}^{r}_{u}$ for users and $\mathbf{H}^{r}_{v}$ for items, which capture complex high-order relationships and dependencies.
(7) $\mathbf{H}^{r}_{u}=Norm(LeakyReLU({x_{u}^{r}}\mathbf{W}_{u}^{r})),$
$\quad\mathbf{H}^{r}_{v}=Norm(LeakyReLU({x_{v}^{r}}\mathbf{W}_{v}^{r})),$
@@
After deriving the user and item hypergraphs for different ratings, we incorporate implicit high-order correlations into the updated representations of users and items. Drawing from traditional hypergraph convolution methods (Feng et al., 2019; Ji et al., 2020), we follow the spectral hypergraph convolution framework. The process begins by aggregating node embeddings to form hyperedge representations via hyperedge convolution. These hyperedge embeddings are then re-aggregated to generate updated node representations. The hypergraph convolution, applied independently to each rating subgraph, is formally defined as follows:
@@
Here, $\mathbf{D}_{u}$ and $\mathbf{D}_{v}$ represent the vertex degree matrices associated with the learned user hypergraph $\mathbf{H}_{u}^{r}$ and item hypergraph $\mathbf{H}_{v}^{r}$, respectively. These matrices play a crucial role in the re-scaling of the resulting embeddings.
@@ 4.4. Attention-Based Cross-View Aggregation
As shown in Eq. (2) and Eq. (8), each node gains a local embedding for the local topology and a global embedding for higher-order dependencies within each rating subgraph. However, downstream tasks require compact embeddings, and traditional pooling methods (e.g., average or sum) often overlook the importance of relations between different node representations.
To overcome this limitation, we propose an attention-based cross-view contrastive aggregation mechanism to reinforce relationships between adjacent ratings. Formally, let $\mathcal{N}_{r}$ denote the two ratings adjacent to rating $r$, and let $x^{r}$ represent the local embeddings of nodes in the $r$-th rating subgraph. The message passing between adjacent ratings is carried out as follows:
(9) $z^{r}=x^{r}+\sum\limits_{r^{\prime}\in\mathcal{N}_{r}}\lambda^{r^{\prime}}\cdot x^{r^{\prime}},$
@@
where $h^{r}$ signifies the global embeddings of nodes within the $r$-th rating subgraph. Similar to $\lambda^{r^{\prime}}$ in Eq. (10), $\delta^{r^{\prime}}$ can be calculated according to the similar computation process like Eq. (10). Subsequently, we consolidate all nodes’ global and local embeddings across rating subgraphs into a unified vector representation using a summation operation. Finally, we further process this intermediate output through a linear operator:
(12) $z[u/v]=tanh(\mathbf{W}_{u/v}^{z}\sum\limits_{r\in{\mathcal{R}}}z^{r}[u/v]),$
$\Gamma[u/v]=tanh(\mathbf{W}_{u/v}^{\Gamma}\sum\limits_{r\in{\mathcal{R}}}\Gamma^{r}[u/v]),$
where $\mathbf{W}_{u/v}^{z}$ and $\mathbf{W}_{u/v}^{\Gamma}$ denote the weight matrices in the nonlinear activation function $tanh(\cdot)$, $z[u/v]$ and $\Gamma[u/v]$ represent local and global embeddings of users/items after cross-view aggregation. The final representation of users and items, $e[u/v]$, is the combination of local embedding and global embedding:
@@
We employ contrastive learning to enhance the complementarity between local collaborative embeddings from user-item interaction graphs and global embeddings from hypergraphs linking users and items. In this framework, different views of the same user or item form positive pairs, while divergent views form negative pairs, providing auxiliary supervision from both local and global spaces. This is formalized with a global-local contrastive loss for user representations, based on the InfoNCE framework, as follows:
(14) $\mathcal{L}_{P}^{\mathbf{U}}=\sum_{u\in\mathbf{U}}-\log\frac{exp((z[u]\cdot\Gamma[u]/\gamma))}{\sum_{u^{\prime}\in\mathbf{U}}exp((z[u]\cdot\Gamma[u^{\prime}]/\gamma))},$
$\mathcal{L}_{P}^{\mathbf{V}}=\sum_{v\in\mathbf{V}}-\log\frac{exp((z[v]\cdot\Gamma[v]/\gamma))}{\sum_{v^{\prime}\in\mathbf{V}}exp((z[v]\cdot\Gamma[v^{\prime}]/\gamma))}.$
The final global-local contrastive loss is the sum of user objective and item objective:
@@ 4.5. Bilinear Decoder and Model Optimization
In line with prior work (Berg et al., 2017; Shen et al., 2021), we adopt a bilinear decoder for the purpose of reconstructing edges within the user-item graph, while considering different ratings as distinct classes. We denote the reconstructed rating matrix as $\check{\mathbf{R}}$, which effectively captures the relationships between users and items. The decoder functions by bi-linearizing the potential rating levels, generating a probability distribution, and subsequently applying a softmax function:
@@
However, traditional CE loss overlooks the sequential and comparable nature of ratings. For example, if the true rating is 3, CE treats predictions like (0.1, 0.15, 0.3, 0.25, 0.2) and (0.05, 0.2, 0.3, 0.35, 0.1) equally, even though the latter is more accurate. To address this, we enhance CE by framing the task as multi-class classification, assigning 1 to the correct rating and a smaller value, $l_{close}$, to adjacent ratings. This adjustment ensures that predictions closer to the true rating receive a lower loss, better capturing both magnitude and order information. By adopting balanced Cross-Entropy (bCE) loss, we effectively integrate the rating’s nuances into our model, improving accuracy while maintaining a balance with $l_{close}$ typically set to 0.1.
@@
We denote the true rating label as $r[i,j]$ and the predicted rating as $\hat{r}[i,j]$ for the pair $(i,j)$. We also introduce hyperparameters $\alpha$ and $\beta$ to optimize the contrastive learning tasks. This results in the final loss function, defined as follows:
@@ 5. Experiment
In this section, we evaluate the performance of our proposed method through extensive experiments and answer the following questions:
@@
  • (RQ2) How does MHCL perform compared to review and rating-based recommendation methods after matrix completion?
  • (RQ3) What are the effects of different modules in MHCL on performance?
  • (RQ4) How does MHCL mitigate long-tail distribution and data sparsity problems?
  • (RQ5) How do different hyperparameter settings affect the performance?
  • (RQ6) How does MHCL perform in terms of efficiency compared to baselines? Experiment results are in Sec. A.6.1 in Supplement.

### v2

@@ 3. Problem Definition
We let $\mathbf{U}=\{u_{1},u_{2},\ldots,u_{M}\}$ denote the set of users and let $\mathbf{V}=\{v_{1},v_{2},\ldots,v_{N}\}$ denote the set of items, where $M$ and $N$ denote the number of users and items, respectively. We typically use a binary matrix $\mathbf{R}$ to store a historical rating from a user to an item and $\mathcal{R}$ is the set of ratings $(\textit{e}.\textit{g}.,\mathcal{R}=\{1,2,3,4,5\})$, where $r_{uv}$ indicates a historical rating from a user $u$ to an item $v$ while $r_{u,v}=0$ means that item $v$ is unexposed to user $u$. Here we represent interaction data as a user-item bipartite graph $G$, and there are only interactions between users and items. The task of matrix completion is to predict missing ratings in the bipartite graph.
As previously discussed, it is noteworthy to highlight that contemporary developments in the realm of GNNs have predominantly seen the emergence of matrix completion techniques within the context of a bipartite user-item graph. This trend underscores the adoption of GNNs to tackle the challenging matrix completion task. In essence, this paradigm shift capitalizes on the inherent power of GNNs to capture complex relationships and dependencies within the user-item interaction data, thereby offering a promising avenue for advancing recommender systems and collaborative filtering approaches. Key notations used in the paper are summarized in Table 1.
@@ 4. Methodology
In this section, we provide a comprehensive overview of the proposed MHCL, as illustrated in Figure 2. MHCL comprises four key components: (1) Subgraph Partition and Graph Convolution: MHCL begins by partitioning the bipartite user-item graph into subgraphs based on different ratings. Graph convolutions are applied to these subgraphs to capture collaborative signals and extract local topological features for users and items across various rating contexts. (2) Multi-Channel Cross-Rating Contrastive Learning: To enhance interaction between adjacent rating subgraphs, MHCL employs a multi-channel cross-rating contrastive learning module. This module promotes information exchange between rating channels, addressing data sparsity and mitigating the long-tail distribution problem. (3) Adaptive Hypergraph Structure Learning: Using local embeddings from the subgraphs, MHCL integrates an adaptive hypergraph structure learning module. This module constructs hypergraphs by identifying high-order dependencies among nodes of the same type (e.g., users with users, items with items), capturing global topological features unique to each rating. This process helps mitigate over-smoothing and enhances the model’s ability to learn higher-order correlations. (4) Attention-Based Cross-View Aggregation: MHCL aggregates local and global representations through an attention-based cross-view aggregation module. By maximizing the consistency between embeddings from diverse views, it leverages global-local contrastive learning techniques to obtain robust final representations.
Through these four components, MHCL delivers a comprehensive solution for matrix completion and recommendation tasks, leveraging collaborative signals, topological features, and higher-order dependencies to enhance the accuracy and efficiency of our model.
@@ 4.1. Subgraph Partition and Graph Convolution
(1) $x^{r}_{0}[i]=Concat(\mathbf{E}^{r}[\mathbf{id}[i]],\mathbf{E}^{r}[\mathbf{ra}[i]],\mathbf{E}^{r}[\mathbf{oh}[i]]),$
where $\mathbf{E}^{r}\in\mathbb{R}^{(M+N+3)\times d}$ denotes $r$-th trainable parameter matrix in $\mathbf{E}\in\mathbb{R}^{\mathcal{|R|}\times(M+N+3)\times d}$, and $x^{r}_{0}[i]$ denotes the initializing embedding of the $i$-th node in the $r$-th rating subgraph. The node embeddings are the input of the message-passing layer.
Then, we systematically partition the graph into distinct rating subgraphs based on the different types of rating edges. For example, with five rating types, we create five rating subgraphs, each containing node embeddings specific to its rating category. MHCL then operates independently on these subgraphs, applying graph convolutions to derive node embeddings within each subgraph. Unlike conventional GCNs, MHCL forgoes the use of a learnable matrix for feature transformation, opting for an explicit propagation process, which can be expressed as follows:
@@ 4.2. Multi-Channel Cross-Rating Contrastive Learning
In contrast to traditional heterogeneous graphs that often feature multiple edge types (e.g., click, purchase, add to cart), we encounter a unique scenario in our context where different types of ratings are inherently comparable. Typically, adjacent ratings indicate that a user’s preference for an item is relatively similar, leading to correspondingly similar node representations. However, the presence of imbalanced rating distributions, stemming from the long-tail distribution inherent in real datasets, adds complexity to the situation. A critical challenge arises due to our proposed MHCL of partitioning the heterogeneous graph into distinct subgraphs based on different rating types, which inadvertently results in the loss of both magnitude and order information between the various ratings when performing graph convolution.
To address this challenge, we introduce an innovative multi-channel cross-rating contrastive learning module. We treat diverse rating subgraphs as distinct (rating) channels, each offering a unique perspective on the data, and we leverage cross-view contrastive learning to capture cooperative associations. Furthermore, we introduce mutual enhancement mechanisms to align adjacent channels, leveraging their inherent correlations to strengthen representations and bolster the effectiveness of MHCL. This module extracts richer learning signals, improving the representations of less frequent ratings and mitigating rating data sparsity. Specifically, we employ the well-established InfoNCE loss (Gutmann and Hyvärinen, 2010) built upon cross-rating representations of users and items. Here, we consider a user’s embedding learned from the current rating subgraph and the user embedding from the next adjacent rating subgraph as positive pairs, respecting the order of ratings. The mathematical equations are presented as follows:
(3) ${\mathcal{L}_{S(k)}^{\mathbf{U}}=\sum_{u\in\mathbf{U}}-\log\frac{exp((x_{u}^{(k)}\cdot x_{u}^{(k+1)}/\tau))}{\sum_{u^{\prime}\in\mathbf{U-u}}exp((x_{u}^{(k)}\cdot x_{u^{\prime}}^{(k+1)}/\tau))},}$
where $x_{u}^{(r)}$ is the normalized output embedding of rating subgraph $r$, $\tau$ is the temperature hyper-parameter of softmax, $u^{\prime}$ represents a node distinct from node $u$.
@@ 4.3. Adaptive Hypergraph Structure Learning
As previously discussed, the embedding derived from Eq. (2) effectively encapsulates collaborative signals and topological information about users and items within distinct rating subgraphs. It is worth noting that the inherent topological similarities among nodes often mirror implicit similarities in user preferences or item features. In light of this insight, MHCL endeavors to transcend the conventional limitation inherent in bipartite graphs where edges typically link nodes of the same type, and involves the hypergraph structure learning for both users and items. Through hypergraph convolution, MHCL adaptively captures global information that transcends the boundaries of node types. Drawing inspiration from prior methodologies (Wei et al., 2022; Xia et al., 2022), we employ the MLP in the computation of hyperedge assignments for users and items within each rating subgraph. This process yields hypergraphs $\mathbf{H}^{r}_{u}$ for users and $\mathbf{H}^{r}_{v}$ for items, each reflecting the intricate relationships and dependencies that exist within the respective rating contexts:
(7) ${\mathbf{H}^{r}_{u}}=normalize(LeakyReLU({x_{u}^{r}}\mathbf{W}_{u}^{r})),\quad{\mathbf{H}^{r}_{v}}=normalize(LeakyReLU({x_{v}^{r}}\mathbf{W}_{v}^{r})),$
@@
Upon deriving the user hypergraph and item hypergraph associated with various ratings, we possess the means to incorporate implicit high-order correlations among nodes into the updated representations of both users and items. Drawing inspiration from conventional hypergraph convolution techniques (Feng et al., 2019; Ji et al., 2020), we adhere to the established framework of spectral hypergraph convolution for its balance between computational efficiency and representational power, which aligns well with our needs. The process unfolds as follows: Initially, node embeddings are amalgamated to form hyperedge representations through hyperedge convolution. Subsequently, these hyperedge embeddings are re-aggregated to create the novel node representations. The hypergraph convolution, which operates independently on each rating subgraph, is formally defined as follows:
@@
Here, $\mathbf{D}_{u}$ and $\mathbf{D}_{v}$ represent the vertex degree matrices associated with the learned user hypergraph $\mathbf{H}_{u}^{r}$ and item hypergraph $\mathbf{H}_{v}^{r}$, respectively. These matrices play a crucial role in the rescaling of the resulting embeddings.
@@ 4.4. Attention-based Cross-View Aggregation
As elucidated in Eq. (2) and Eq. (9), each node acquires a local embedding that captures the intricacies of the local topology, and a global embedding that encapsulates higher-order dependencies within each rating subgraph. However, in the context of downstream tasks, the need often arises for a more compact node embedding. Traditional methods commonly resort to straightforward pooling operations (such as average or sum pooling) applied to embeddings from various perspectives to obtain condensed representations. However, these operations tend to overlook the significance of node representations corresponding to diverse relations.
To address this limitation, we introduce an attention-based cross-view contrastive aggregation mechanism that aims at strengthening relationships between adjacent ratings. Formally, let $\mathcal{N}_{r}$ represent the two ratings that are adjacent to rating $r$, and let $x^{r}$ denote the local embeddings of nodes within the $r$-th rating subgraph. The process of message passing between adjacent ratings is executed as follows:
(9) $z^{r}=x^{r}+\sum\limits_{r^{\prime}\in\mathcal{N}_{r})}\lambda^{r^{\prime}}\cdot x^{r^{\prime}},$
@@
where $h^{r^{\prime}}$ signifies the global embeddings of nodes within the ${r^{\prime}}$-th rating subgraph. Similar to $\lambda^{r^{\prime}}$ in Eq. (10), $\delta^{r^{\prime}}$ can be calculated according to the similar computation process. Subsequently, we consolidate the global and local embeddings of all nodes across rating subgraphs into a unified vector representation using a summation operation. Finally, we further process this intermediate output through a linear operator:
(12) $z[u/v]=tanh(\mathbf{W}_{u/v}^{z}\sum\limits_{r\in{\mathcal{R}}}z^{r}[u/v]),\quad\Gamma[u/v]=tanh(\mathbf{W}_{u/v}^{\Gamma}\sum\limits_{r\in{\mathcal{R}}}\Gamma^{r}[u/v]),$
where $\mathbf{W}_{u/v}^{z}$ and $\mathbf{W}_{u/v}^{\Gamma}$ denote the weight matrices in the nonlinear activation function $tanh(\cdot)$, $z[u/v]$ and $\Gamma[u/v]$ represent local and global embeddings of users/items after cross-view aggregation. The final representation of users and items, $e[u/v]$, is the combination of local embedding and global embedding of the user/item:
@@
Subsequently, we employ contrastive learning to enhance the complementarity between local collaborative embeddings from user-item interaction graphs and global embeddings from hypergraphs linking users and items. In this framework, different views of the same user or item form positive pairs, while divergent views form negative pairs, providing auxiliary supervision from both local and global spaces. This is formalized with a global-local contrastive loss for user representations, based on the InfoNCE framework, as follows:
(14) $\mathcal{L}_{P}^{\mathbf{U}}=\sum_{u\in\mathbf{U}}-\log\frac{exp((z[u]\cdot\Gamma[u]/\gamma))}{\sum_{u^{\prime}\in\mathbf{U}-u}exp((z[u]\cdot\Gamma[u^{\prime}]/\gamma))},\quad\mathcal{L}_{P}^{\mathbf{V}}=\sum_{v\in\mathbf{V}}-\log\frac{exp((z[v]\cdot\Gamma[v]/\gamma))}{\sum_{v^{\prime}\in\mathbf{V}-v}exp((z[v]\cdot\Gamma[v^{\prime}]/\gamma))}.$
The final global-local contrastive loss is the sum of the user objective and the item objective:
@@ 4.5. Bilinear Decoder and Model Optimization
In line with prior work (Berg et al., 2017; Shen et al., 2021), we adopt a bilinear decoder to reconstruct edges within the user-item graph, while considering different ratings as distinct classes. We denote the reconstructed rating matrix as $\check{M}$, which effectively captures the relationships between users and items. The decoder functions by bilinearizing the potential rating levels, generating a probability distribution, and subsequently applying a softmax function:
@@
However, traditional CE loss overlooks the sequential and comparable nature of ratings. For example, if the true rating is 3, CE treats predictions like (0.1, 0.15, 0.3, 0.25, 0.2) and (0.05, 0.2, 0.3, 0.35, 0.1) equally, even though the latter is more accurate. To address this, we enhance CE by framing the task as multi-class classification, assigning 1 to the correct rating and a smaller value, $l_{close}$, to adjacent ratings as follows:
@@
where we denote the true rating label as $r[i,j]$ and the predicted rating as $\hat{r}[i,j]$ for the pair $(i,j)$. Then we introduce $\alpha$ and $\beta$ to optimize the contrastive learning tasks. This results in the final loss function, defined as follows:
@@ 5. Experiment
In this section, we evaluate the performance of our proposed MHCL through extensive experiments and answer the following questions:
@@
  • (RQ2) How does MHCL perform compared to rating and review-based recommendations after matrix completion?
  • (RQ3) How does MHCL perform on large-scale datasets?
  • (RQ4) What are the effects of different modules in MHCL on performance?
  • (RQ5) How does MHCL mitigate long-tail distribution and data sparsity problems?
  • (RQ6) How do different hyperparameter settings affect the performance?
  • (RQ7) How does MHCL perform in terms of efficiency compared to baselines?
  • (RQ8) What insights can be obtained from case studies of the proposed MHCL?